1. Consider rolling two dice once. Calculate the probability of the...

( a) rolling a sum of 5 or an even sum

(b) rolling a sum of 7 or 6

( c) rolling a sum of 10 or an odd sum

2. Six(6) woman , 4 men,3 boys and 7 girls enter a competition. They each have an equally likely chance of winning. What is the probability that the winner will be...

( a) female

( b) a child

( c) a man

3. A basket contains 12 black marbles ,14 red marbles and 12 green marbles. What is the probability of picking a black marble first and then a red marble with replacement

4. A bag contains 4 green,3 black and 6 orange cards. A card is drawn from the bag and kept, and then another card is drawn. Find the probability that

( a) both cards will be orange

( b) both cards will be black

( c) the first card will be orange and the second will be black

( d) neither will be orange.


1
Expert's answer
2021-06-25T12:39:32-0400

1.

a) 4 / 36 + 1/2 = 0.611

b) 5 / 36 + 6 / 36 = 0.306

c) 3 / 36 + 1 / 2 = 0.583

2.

a) (6+7) / 20 = 0.65

b) (3+7) / 20 = 0.50

c) 4 / 20 = 0.20

3.

(12 / (12+14+12))((14 / (12+14+12)) = 0.116

4.

a) (6 / 13)(5 / 12) = 0.192

b) (3 /13)(2 / 12) = 0.038

c) (6 / 13)(3 / 12) = 0.115

d) (7 / 13)(6 / 12) = 0.269